OpenStudy (anonymous):

Determine whether the sequence is arithmetic or geometric. Sequence 1: –10, 20, – 40, 80, ... Sequence 2: 15, – 5, – 25, – 45, ... Which of the following statements are true regarding Sequence 1 and Sequence 2. A) Sequence 2 is arithmetic and Sequence 1 is geometric. B) Both sequences are arithmetic. C) Both sequences are geometric. D) Sequence 1 is arithmetic and Sequence 2 is geometric.

OpenStudy (abb0t):

@iambatman

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

Haha, well there is a nice way to check this, lets look at the first sequence \[-10, 20, -40, 80, ...\] do you see a pattern here?

OpenStudy (abb0t):

yes I do.

OpenStudy (anonymous):

Ok well lets look at the first two terms, -10, 20 notice if you multiply -10 by -2 you will get 20? Would the same work if you multiply 20 by -2 to get the next term (-40)?

OpenStudy (anonymous):

Basically it's an arithmetic sequence if \[t_2-t_1 = t_3 - t_1\]and it's a geometric sequence if \[\frac{ t_2 }{ t_1 } = \frac{ t_3 }{ t_2 }\] that gives you your common ratio. Where the t represents the term, and the subscript is the address of the term, so \[t_1 = -10~~~t_2 = 20\] etc

OpenStudy (triciaal):

an arithmetic sequence has the same difference between the terms and the geometric sequence has the same ratio between the terms.

OpenStudy (texaschic101):

in an arithmetic sequence, the next number is found by multiplying...and each term has to be multiplied by the same number. In a geometric sequence, the next number is found by dividing...and each term has to be divided by the same number.